Opened 8 years ago

Last modified 7 years ago

#422 new flight processing

BB11_01, flight day 272a/2011, Brecon Beacons

Reported by: knpa Owned by: jaho
Priority: alpha 4 high Milestone: 2011 data processing completion
Component: Processing: general Keywords:
Cc: Other processors: adbe,knpa

Description (last modified by adbe)

Data location: ~arsf/arsf_data/2011/flight_data/uk/BB11_01-2011_272a_Brecon_Beacons

Data arrived from ARSF via network transfer on 06/10/2011

BB11_01 Scientific objective:

BB11_01 Priority: a4h

PI: Shaun Lewis

Sensors:

Camera (23/11/2011)
Eagle (23/11/2011)
Hawk (23/11/2011)
LiDAR (23/11/2011)

Change History (20)

comment:1 Changed 8 years ago by adbe

Started to process navigation

comment:2 Changed 8 years ago by adbe

Split the project into 272a and 272c, for navigation reasons

comment:3 Changed 8 years ago by adbe

Started processing hyperspectral

comment:4 Changed 8 years ago by jaho

Started LiDAR processing.

comment:5 Changed 8 years ago by jaho

LiDAR ready for delivery check.

comment:6 Changed 8 years ago by adbe

Hyperspectral ready for delivery check

comment:7 Changed 8 years ago by emca

Delivery checked LiDAR data - no issues. Ready to deliver.

comment:8 Changed 8 years ago by emca

Processing camera - removed images 1-19 as they are not part of the study area.

comment:9 Changed 8 years ago by emca

Camera processing complete - ready for delivery check

comment:10 Changed 8 years ago by jaho

Hyperspectral delivery check complete.

There's no flight letter on the first page of the ReadMe. Everything else looks good.

comment:11 Changed 8 years ago by adbe

Added flight letter onto front page of hyperspectal ReadMe

comment:12 Changed 8 years ago by knpa

Gary says that the PI should be able to get their own vectors via PGA.
Do not wait for vectors to make delivery but deliver with caveat that we have not made this usual check.

comment:13 Changed 8 years ago by jaho

Camera delivery check.

KML file should be named BB11_01-2011_272a.kml. Everything else looks good.

comment:14 Changed 8 years ago by adbe

Camera, Eagle, Hawk and LiDAR on disk 117 and ready to deliver on 08/11/2011 pending payment

comment:15 Changed 8 years ago by emca

Reprocessed LiDAR data for pitch correction, delivery folder BB11_01-272a-lidar-20111122.

Will need to be delivery checked again, and rewriting to disk 117 once delivery check is complete.

Should remove delivery folder BB11_01-272a-lidar-20111013 once we are happy with the reprocessed data.

comment:16 Changed 8 years ago by adbe

LiDAR delivery check:

No problems, ready to deliver

comment:17 Changed 8 years ago by jaho

Updated LiDAR ReadMe to mention vectors.

comment:18 Changed 8 years ago by adbe

  • Description modified (diff)

Camera, Hyperspectral and LiDAR delivered to Shaun Lewis on 23/11/2011 on disk 117

comment:19 Changed 7 years ago by mark1

Hyperspectral hawk bad pixel files and XML files have been re-created and re-delivered to PI (via email). This corrects sample number in bad pixel text files and Eagle IFOV value in XML and extra comments included.

comment:20 Changed 7 years ago by adbe

Hyperspectral sct's

lineeaglehawk
1-0.05-0.05
2-0.08-0.08
3-0.08-0.05
4-0.09-0.12
5-0.03-0.03
6-0.10-0.12
7-0.05-0.08
Note: See TracTickets for help on using tickets.